OpenStudy (anonymous):

please help I don't get it? 60% of a number is 300. What is 25% of the number

OpenStudy (anonymous):

@igreen

OpenStudy (anonymous):

In percentage problems the word "is" means = and the word "of" means times. Therefore your problem would look like this: 60 = x(300) now solve for x by dividing both sides by 300 60 / 300 = x now divide this out to a decimal or get a denominator of 100 which would be the percentage. .20 = 20% or 20 / 100 = ? That is the way I try to teach percentage problems. Therefore this is your answer → 20 %

OpenStudy (anonymous):

Then 25% of 20 is 5! @amir7816

OpenStudy (anonymous):

actually Amycarter in percentage "is" means equals, and "of" means times

OpenStudy (anonymous):

nvm you had it I read it wrong